Activity during this reporting period has been at a low level. The seismic records were dominated by small rockfall and pyroclastic flow signals from the unstable part of the dome above the Galway's Wall on the southern side of the volcano. A slight increase in the number of long-period earthquakes occurred in the middle of the reporting period. EDM and GPS measurements around the volcano indicate that ground deformation is still very slow and confined to areas close to English's Crater.
Early in the week views were obtained of the scar left by the collapse on April 11th. It did not appear to be much larger than the Easter scar, however the collapses had eaten back a further 100 m into the pre-September material on the south-west side of the dome. The activity was concentrated in this region with continued growth of the lobe rapidly replacing the material lost on the 11th. Only very small rockfalls were noted from the northern or eastern flanks of the dome.
The head wall of the lobe continued to move south towards the remains of the Galway's Wall. A GPS/laser binocular survey conducted on the 15th showed that the southern summit continues to decrease in altitude, but at a slower rate. The southern summit was measured at 947 m a.s.l. and no change was detected for the northern summit. The most noticeable change in the southern region was the increase in the width of the lobe, to about 180 m. Most of this expansion had occurred on its west side, and fresh material is now encroaching on the remains of the September scar. To the east, the edge of the lobe is now against the Easter collapse scar and the small rockfalls observed going towards the south-east crater wall have been from this area. During the week, the smooth surface of the lobe became increasingly fractured. The area around some of these fractures were stained brown by gas emissions and the saddle between the summits also showed periods of vigorous degassing and ashy steam production.
The chute over the Galway's Wall is continuing to widen and fill slowly, however the material is still at about 50( to the horizontal. There is some evidence to suggest that the pyroclastic flows on the 11th eroded material from the base of the chute and this material is now being replaced by the short-runout rockfalls and small pyroclastic flows. The soufriere area continues to be filled by pyroclastic flow deposits, but flows did not progress much further than this region during this period.
Observations were made from a new observation post on the South Soufriere Hills. This allows unprecedented views of the Galway's Wall area and the rockfall/pyroclastic flow activity.
The level of seismic activity was quite low during the week, with a continuation of rockfalls and small pyroclastic flows to the south of the dome. There were few earthquakes, with only one hybrid earthquake swarm (Table 2). The number of long-period earthquakes increased mid-week, although the number that triggered rockfalls remained low.
Table 1: Earthquake types
These earthquake counts are of events that triggered the broadband network event recording systems between 0000 and 2400 each day (local time).
Date VT LP Hybrid Dome RF LPRF* 13 April 96 0 2 13 26 2 14 April 96 1 3 11 20 3 15 April 96 5 4 77 30 3 16 April 96 1 8 10 27 6 17 April 96 0 25 1 59 7 18 April 96 1 29 4 34 5 19 April 96 0 12 2 16 2
*LPRF: LP earthquake followed by rockfall signal. The LPs and rockfalls in an LPRF signal are also counted in their respective columns.
An EDM occupation of the northern triangle (Windy Hill - Farrell's -St. George's Hill) on April 15th was incomplete due to poor atmospheric conditions preventing a return signal reaching the instrument on the St. George's Hill - Farrell's leg of the occupation. One result of this occupation was a 40 millimetre shortening of the Windy Hill - Farrell's slant distance with respect to the previous occupation on April 2nd. This change was almost twice as large as any change seen on this between successive occupation before. The northern triangle has always been the subject of wide fluctuations attributed to atmospheric conditions and the long lines involved but this change seemed out of place.
A subsequent full occupation of the northern triangle on April 18th indicated that the Windy Hill - Farrell's slant distance had increased by 22 millimetres with respect to the previous occupation on April 15th. It therefore seems likely that the large change observed on April 15th can be attributed to atmospheric effects but further occupations in conjunction with GPS surveys to the same points will be carried out to check the changes.
A GPS occupation of EASTNET (Long Ground-Whites-M18 Harris Lookout-Windy Hill-Farrell's) was carried out on April 16th. No significant changes were observed for the positions of all the stations except Farrell's. In view of the large change on the Windy Hill-Farrell's EDM line on April 15th a long occupation GPS survey between Harris Lookout (M18) & Farrell's was performed on April 17th. Both the GPS and EDM techniques are in agreement that there is slow movement of the Farrell's site away from the crater.
Dome Volume Measurements
An estimation has been made of the volume of the lobe over the Galway's Wall from the GPS/laser binocular data obtained on the 15th. The data constrain the maximum size of the lobe and suggests that it is around 0.9x106 cubic metres.
The MiniCOSPEC machine is currently in Arizona with Billy Darroux of the MVO. This is for a workshop on the monitoring of sulphur dioxide at active volcanoes using COSPEC. The workshop lasts for one week and it is planned that all COSPEC machines will be calibrated to a common standard.
No water or ash samples were collected during this reporting period.
Dust sampling in Plymouth and its surroundings continues with pumps at Cork Hill school, the Port, Kinsale checkpoint, Monlec and various other sites. The samples will be sent to the UK for analysis.
Dr Paul Cole (University of Luton, UK)